jeremy pays for lunch using 75 coins consisting of quarters and nickles. If the total for his lunch is $13.55, find the number of quarters and nickles.
Answer by oberobic(2304)

n = number of nickels
5n = value of nickels in cents
q = number of quarters
25q = value of quarters in cents
.
5n + 25q = 1355 cents
.
n + q = 75 coins
.
multiply the 2nd equation by 5 and subtract from the first
.
5n + 25q = 1355
5n + 5q = 375
-----------------
20q = 980
q = 49
.
n + q = 75
n = 75 -49
n = 26
.
Check the values
.
25q = 25(49) = 1225 cents
5n = 5(26) = 130 cents
1225 + 130 = 1355 cents
Correct.
.
Answer: 49 quarters and 26 nickels.
.
Done.